# Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities

### v2

@@ II Comparison of fitting quality
Figure S2: Comparison of fitting quality with Eq. S1 and Eq. S2 in the AFMXY phase at $\Delta=0$ and $q=0.3$. Left panel: $S^{(2)}_{A}-0.1338L$ versus $\ln(L)$ with area law coefficient obtained from fitting to Eq. S1. Right panel: $S^{(2)}_{A}-0.1504L$ versus $1/L$ with area law coefficient obtained from fitting to Eq. S2. The chi-squared values of two fittings are also printed on the figure.
Figure S3: Comparison of fitting quality with Eq. S1 and Eq. S2 at the DQCPs of easy-plane $JQ_{3}$ model. The red and blue data points represent the finite-size EE data with an area law term subtracted. The area law term is fitted from Eq. S1 for red data points and Eq. S2 for blue data points.
@@
$\chi^{2}=\sum_{i=1}^{N}\frac{\left(f(x_{i})-y_{i}\right)^{2}}{\sigma_{i}^{2}},$ (S3)
where $\sigma_{i}$ is the uncertainty of data $y_{i}$. $k=N-r$ is the effective number of degrees of freedom, where $N$ is the number of data points and $r$ is the number of fitting parameters. $\chi^{2}/k$ is expected to be close to 1 for a good fit. As shown in Fig. S2, in the AFMXY phase, the fitting quality is compared with Eq. S1 and Eq. S2 and the $\chi^{2}/k$ values are 0.395 and 4.185 separately. $\chi^{2}/k$ should be typically distributed within the range $[1-\sqrt{2/k},1+\sqrt{2/k}]$ where for this case the range is $[0.465,1.535]$. In our simulation, the precise estimation of statistic errors can be affected by limited number of Monte Carlo bins so that we attribute the small deviation of 0.395 to 0.465 to problematic errorbars and 4.185 is apparently far away from the ideal range. We thus conclude that Eq. S1 fits the data better than Eq. S2.
Similarly, for DQCPs of easy-plane $JQ_{3}$ model, the comparison of fitting quality with Eq. S1 and Eq. S2 is also performed and the results are shown in Fig. S3. By comparing their $\chi^{2}/k$ values, we conclude that Eq. S1 in all cases fit better than Eq. S2. Note that for $\Delta=1.0$ which is the standard $JQ_{3}$ model, it has already been shown in previous study Song et al. (2023a) that Eq. S1 fits the better. In that case we only present the comparison for $\Delta<1.0$ in this manuscript.

### v3

@@ Appendix B Comparison of fitting quality
Figure 7: Comparison of fitting quality with Eq. 10 and Eq. 11 in the AFMXY phase at $\Delta=0$ and $q=0.3$. Left panel: $S^{(2)}_{A}-0.1338L$ versus $\ln(L)$ with area law coefficient obtained from fitting to Eq. 10. Right panel: $S^{(2)}_{A}-0.1504L$ versus $1/L$ with area law coefficient obtained from fitting to Eq. 11. The chi-squared values of two fittings are also printed on the figure.
Figure 8: Comparison of fitting quality with Eq. 10 and Eq. 11 at the DQCPs of easy-plane $JQ_{3}$ model. The red and blue data points represent the finite-size EE data with an area law term subtracted. The area law term is fitted from Eq. 10 for red data points and Eq. 11 for blue data points.
@@
$\chi^{2}=\sum_{i=1}^{N}\frac{\left(f(x_{i})-y_{i}\right)^{2}}{\sigma_{i}^{2}},$ (12)
where $\sigma_{i}$ is the uncertainty of data $y_{i}$. k=Nrk=N-r is the effective number of degrees of freedom, where $N$ is the number of data points and $r$ is the number of fitting parameters. $\chi^{2}/k$ is expected to be close to 1 for a good fit. As shown in Fig. 7, in the AFMXY phase, the fitting quality is compared with Eq. 10 and Eq. 11 and the $\chi^{2}/k$ values are 0.395 and 4.185 separately. $\chi^{2}/k$ should be typically distributed within the range $1-\sqrt{2/k},1+\sqrt{2/k}]$ where for this case the range is $[0.465,1.535]$. In our simulation, the precise estimation of statistic errors can be affected by limited number of Monte Carlo bins so that we attribute the small deviation of 0.395 to 0.465 to problematic errorbars and 4.185 is apparently far away from the ideal range. We thus conclude that Eq. 10 fits the data better than Eq. 11.
Similarly, for DQCPs of easy-plane $JQ_{3}$ model, the comparison of fitting quality with Eq. 10 and Eq. 11 is also performed and the results are shown in Fig. 8. By comparing their $\chi^{2}/k$ values, we conclude that Eq. 10 in all cases fit better than Eq. 11. Note that for $\Delta=1.0$ which is the standard $JQ_{3}$ model, it has already been shown in previous study [13] that Eq. 10 fits the better. In that case we only present the comparison for $\Delta<1.0$ in this manuscript.